Multiply:
(x^2+x+4)(x-4)
Add:
(8x^2-xy+y^2)+(-x^2-4xy+6y^2)
Factor By Grouping:
4x^3-28x^2-3x+21
Factor:
3c^7-192c^4

I have been working on these problems for quite sometime.. Please help, if possible, show all work so I can understand better! Thank you!

Found 2 solutions by stanbon, checkley77:
Answer by stanbon(57395)

Multiply:
(x^2+x+4)(x-4)
---
= (x^2+x+4)x - (x^2+x+4)4
-----
= x^3+x^2+4x -4x^2-4x-16
---
= x^3-3x^2+16
----------------------
Add:
(8x^2-xy+y^2)+(-x^2-4xy+6y^2
---
= 7x^2-5xy+7y^2
------------------------
Factor By Grouping:
4x^3-28x^2-3x+21
---
= 4x^2(x-7)-3(x-7)
---
= (x-7)(4x^2-3)
------------------------
Factor:
3c^7-192c^4
---
= 3c^4[c^3-64]
---
= 3c^4(c-8)(c^2+8c+16)
=========================
Cheers,
Stan H.

Answer by checkley77(12569)

Multiply:
(x^2+x+4)
(x-4)
---------------------
X^3+X^2+4X-4X^2-4X-16
X^3-3X^2-16 ans.
Add:
(8x^2-xy+y^2)
(-x^2-4xy+6y^2)
--------------------
7x^2-5xy+7y^2 ans.
Factor By Grouping:
4x^3-28x^2-3x+21
4x^2(x-7)-3(x-7)
(4x^2-3)(x-7) ans.
Factor:
3c^7-192c^4
3c^4(c^3-64)
3c^4(c-4)(c^2+4c+16) ans.